find the equation of a line with a y-intercept of 3 and goes through (-2,1)
Answer by algebrahouse.com(1659)

Since the y-intercept is 3, then (0,3) is another point on the line.

y - y1
------ = slope
x = x1

(0,3) and (-2,1) are two points on the line

1 - 3
------ = slope
-2 - 0

-2
--- = slope
-2

slope = 1

Point is (-2,1) and slope is 1


Point-slope form for the equation of a line is y - y1 = m(x - x1
(x1,y1) are the coordinates of the point
m is the slope


y - 1 = 1[x - (-2)] {substituted point and slope into point-slope form}
y - 1 = x + 2 {simplified}
y = x + 3 {added 1 to each side}

For more help from me, visit:
www.algebrahouse.com